# Modality Unified Attack for Omni-Modality Person Re-Identification

### v1

@@ I Introduction
Person re-identification (re-id), which seeks to retrieve a specific individual in video surveillance, has drawn significant attention in recent years [1, 2]. Over the past decades, numerous single-modality re-id methods have been proposed for RGB image matching [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], showcasing promising results. However, these RGB-based methods often struggle in low-light or dark environments. To overcome these limitations, the cross-modality re-id models have been introduced [14, 15, 16, 17, 18, 19], which aim retrieving the same person across RGB and near-infrared (NI) modality images. More recently, researchers have been developing multi-modality re-id methods to leverage complementary information from RGB, NI, and thermal infrared (TI) images to create even more robust and effective re-id systems [20, 21, 22, 23].
@@ III-A Problem Definition
Algorithm 1 Modality Unified Attack
2:  for $i$=0 to $\mathcal{I}$ do
3:     Sample batch images $x_{h}$ and generate AEs $x^{\prime}_{h}$ by Eq.2
4:     Feed $x_{h}$ and $x^{\prime}_{h}$ to $\mathcal{M}_{M}$
5:     Calculate Metric Disruption loss $\mathcal{L}^{h}_{MD}$ by Eq.7
6:     Cross-inputs the R/N modalities to N/R modality subnetworks
12:  end for
@@ IV-A Experiments Setup
$aAP=\ \frac{\sum_{i=0}^{N}{mAP}_{i}}{N},$ (16)
$mDR=\frac{aAP-aAP^{\prime}}{aAP},$ (17)
where $aAP^{\prime}$ is evaluated on generated adversarial examples.
@@ IV-C Ablation Studies
V Conclusion
@@
In this paper, we propose the first modality unified attack for omni-modality (single-, cross-, multi-modality) re-id models. Cross Modality Simulated Disruption method is introduced to attack the shared features of the same modality in omni-modality models by simulating features embeddings across different types of re-id models. Multi Modality Collaborative Disruption approach is designed for the attacker to use multi modality features to collaboratively measure the feature disruption and comprehensively damage the information content of the person image. Extensive experiments on various single-, cross- and multi-modality black-box models show that our method can effectively attack the omni-modality re-id models.

### v2

@@ I Introduction
Person re-identification (re-id), which seeks to retrieve a specific individual in video surveillance, has drawn significant attention in recent years [1, 2]. Over the past decades, numerous single-modality re-id methods have been proposed for RGB image matching [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], showcasing promising results. However, these RGB-based methods often struggle in low-light or dark environments. To overcome these limitations, the cross-modality re-id models have been introduced [15, 16, 17, 18, 19, 20], which aim retrieving the same person across RGB and near-infrared (NI) modality images. More recently, researchers have been developing multi-modality re-id methods to leverage complementary information from RGB, NI, and thermal infrared (TI) images to create even more robust and effective re-id systems [21, 22, 23, 24].
@@ III-F Overall Objective
Algorithm 1 Modality Unified Attack
2:  for $i$=0 to $\mathcal{I}$ do
3:     Sample batch images $x_{h}$ and generate AEs $x^{\prime}_{h}$ by Eq.2
4:     Feed $x_{h}$ and $x^{\prime}_{h}$ to $\mathcal{M}_{M}$
5:     Calculate Metric Disruption loss $\mathcal{L}^{h}_{MD}$ by Eq.7
6:     Cross-inputs the R/N modalities to N/R modality subnetworks
12:  end for
@@ IV-A Experiments Setup
$aAP=\ \frac{\sum_{i=0}^{N}{mAP}_{i}}{N},$ (16)
$mDR=\frac{aAP-aAP^{\prime}}{aAP},$ (17)
where $aAP^{\prime}$ is evaluated on generated adversarial examples.
@@ IV-C Ablation Studies
Figure 4: Attention maps of RGB, NI, and TI modalities for both benign and adversarial examples across different re-ID models. The original attention maps correspond to benign images, while the others visualize adversarial examples generated under different loss constraints (MD, MMCD, CMSD). Transreid [12], MMN [65] and EDITOR [23] are token as the single-, cross- and multi-modality re-id models for Visualizations.
To further validate the effectiveness of our proposed methods, we visualize the attention maps for both benign and adversarial examples of RGB, NI, and TI modalities on different types of re-id models. The results in Fig. 4(a) show the attention maps of RGB on single-modality re-id model, which illustrates the MMCD and MMCD can eliminate the attention and CMSD guided different attention attack on the same images. For attention maps of cross-modality re-id model in Fig. 4(b), the MMCD and CMSD continuously disrupt the attentions of target person on both RGB and NI images. There are also obviously comprehensive disruptions of three modality images on multi-modality re-id with our methods, shown in Fig. 4(c). In summary, the visualization results demonstrate that our method effectively and comprehensively perturbs feature representations across different re-id tasks (single-modality, cross-modality, and multi-modality) and modalities (RGB, NI, TI), thereby validating its strong adversarial transferability and attack efficacy.
Figure 5: Analysis of mDR under different perturbation strength and balance weights. The mDR values in the balance weights analysis represent the average across four retrieval settings.
TABLE III: Attack effectiveness against defense methods.
Re-id models Method Randomization JPEG $aAP$ $mDR$
Single-modality None 84.6 83.8 84.2 -
($R^{\prime}$-$R$) MUA 45.2 54.1 49.7 41.0
Cross-modality None 59.6 58.4 59.0 -
($R^{\prime}$-$N$) $MUA$ 49.2 47.4 48.3 18.1
Cross-modality None 54.3 52.2 53.3 -
($N^{\prime}$-$R$) $MUA$ 32.8 31.7 32.3 39.5
Multi-modality None 60.4 59.5 60.0 -
($R^{\prime}N^{\prime}T^{\prime}$-$RNT$) $MUA$ 39.1 36.9 38.0 36.7
Analysis of key parameters. We conduct analysis on the balance weights $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ and different perturbation strength $\epsilon$. To balance the influence of different loss terms in Eq. 14, we employ these $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ three weighting coefficients. We investigate their individual impacts through an ablation study, where we systematically vary one hyperparameter while keeping the others fixed. Fig. 5(a) illustrates the impact of balance weights, showing that high $\lambda_{1}$ and $\lambda_{2}$ values lead to notable mDR declines, potentially forcing overfitting to surrogate model feature disruption. In contrast, performance remains stable across varying $\lambda_{3}$ values, demonstrating the stabilizing effect of the MMCD loss. On top of the experiments, we set $\lambda_{1}$=50, $\lambda_{2}$=50, and $\lambda_{3}$=10, which achieves the best performance. For perturbation strength, larger values improve attack performance, but we chose 8/255 to balance imperceptible perturbations.
IV-D Attack Effectiveness against Defense Method
Tab. III provides supplementary experimental results analyzing the robustness of our approach when confronted with different defense mechanisms. These include input transformation defenses such as JPEG compression [75] with a quality factor set to 60%, as well as noise-based mitigation approaches like Randomization [76]. The experiments were conducted using the single-modality BOT [54] model, cross-modality MMN [65] and multi-modality EDITOR [23] model as the target system for evaluation. It is notable that our method also demonstrates attack effectiveness against these defenses, with 41.0%, 18.1%, 39.5% and 36.7% mDR for omni-modality re-id models.
Figure 6: The rank-5 retrieval results of single-. cross- and multi-modality models before and after our attack. Green boxes denote correctly matched images, red boxes indicate mismatched images, and the first row represents the benign results and the second row represents the adversarial results.
IV-E Visualization
Figure 7: Visualization of perturbations (Pert) and adversarial examples (AE) generated by our MUA framework across three modalities: RGB (top), NI (middle), and TI (bottom).
We visualize the adversarial examples and perturbations generate by our MUA across different modality images. As Fig. 7 shows, the perturbations on RGB images exhibit semantically meaningful patterns correlated with target pedestrians and texture perturbations, while those on NI and TI modalities primarily focus only on texture perturbations. This observation aligns well with our expectations for perturbations on different modalities: RGB images are more susceptible to semantic-level perturbations due to their rich visual information and texture perturbations can also attack them in cross-modality tasks, whereas NI and TI modalities, lacking explicit semantic features, achieve better adversarial transferability through texture-based perturbations that preserve modality characteristics while effectively deceiving models. Furthermore, the adversarial examples across all modalities demonstrate imperceptible perturbations. Our modality-specific generators effectively produce perturbations that are carefully adapted to each modality’s unique imaging characteristics.
We also provide visualization of attack results by showcasing the Top-5 retrieval results from the different types target re-id models (i.e., Transreid [12] (single-modality), MMN [65] (cross-modality) and EDITOR [23] (multi-modality)). Fig. 6 demonstrates that our method successfully launches effective attacks against various model architectures, significantly compromising their retrieval performance.
V Discussion
Impact. The proposed adversarial attack method highlights potential security risks in surveillance systems by demonstrating effective attacks against different types of person re-identification models. These findings expose vulnerabilities that could be exploited by malicious actors, emphasizing the need for more robust defenses. At the same time, the method provides a valuable benchmark for evaluating and improving model robustness. In future work, we will explore how the generated adversarial examples can be used to develop more resilient re-identification systems through adversarial training and robustness testing.
Limitation and future work. The proposed MUA method employs three modality-specific adversarial generators to enable comprehensive omni-modality re-id attacks. While effective, this multi-generator architecture introduces implementation complexity and parameter redundancy. To address these limitations, our future work will develop a diffusion-based unified framework that simultaneously generates effective adversarial samples across all modalities and preserves the distinctive imaging characteristics of each modality.
VI Conclusion
@@
In this paper, we propose the first modality unified attack for omni-modality (single-, cross-, multi-modality) re-id models. Cross Modality Simulated Disruption method is introduced to attack the shared features of the same modality in omni-modality models by simulating features embeddings across different types of re-id models. Multi Modality Collaborative Disruption approach is designed for the attacker to use multi modality features to collaboratively measure the feature disruption and comprehensively damage the information content of perdestrain images. Extensive experiments on various single-, cross- and multi-modality black-box models show that our method can effectively attack the omni-modality re-id models.